OpenStudy (anonymous):

In scientific notation, 8,700 is written: 8.7 x (10^1) 8.7 x (10^2) 8.7 x (10^3) 8.7 x (10^4)

OpenStudy (whpalmer4):

8700 = 8.7 * 10, 8.7*100, 8.7*1000, or 8.7*10000? 10^1= 10 10^2=100 10^3=1000 10^4=10000

OpenStudy (anonymous):

i mean It is: 8.7*103

OpenStudy (anonymous):

10^3 you mean also

OpenStudy (whpalmer4):

Maybe even \(8.7*10^3\) :-)